Phenytoin (Dilantin) 5 mg/kg is ordered for a 40-pound child. It is to be administered in...

  1. Phenytoin (Dilantin) 5 mg/kg is ordered for a 40-pound child. It is to be administered in three equal doses. The drug is available in an oral suspension containing 125 mg/ml. How many mL should be administered per dose?
  1. The order is for 1.2 million units of penicillin G (Bicillin) IM. Available is 600,000 units/ml. How much should the nurse administer?
  2. Order is for 2500 ml D5W over 24 hours. Drop factor is 15 gtt/ml. At what rate do you run the IV?

  1. Order is for 2000 ml D5W over 24 hours. Drop factor is 60 gtt/ml. What rate do you run the IV?
  1. Enoxaparin sodium (Lovenox) 25 mg SC q 12 hours is ordered. The label reads 30 mg/0.3 ml. How much should the nurse administer
  1. The order is for meperidine 50 mg IM q 4 hours prn. The label reads meperidine 75 mg/ml. How much should the nurse administer?
  1. An adult is on continuous IV heparin therapy for thrombophlebitis. The IV contains 15,000 units of heparin in 500 ml of 5% dextrose (D5W) at the rate of 20 ml per hr. How many units per hours is the client receiving?
  1. The order is for Ancef 1 gram IV in 50 cc 5% dextrose (D5W) to run in over 30 minutes every 6 hours. The administration set delivers 10 gtts/cc. What should the drip rate be?
  1. The HCP prescribes haloperidol 4 mg tid oral liquid for the patient diagnosed with schizophrenia. How many milliliters should the nurse administer from a bottle labeled “haloperidol 2 mg/ml,” for each dose?
Answer #1
  1. Phenytoin (Dilantin) 5 mg/kg is ordered for a 40-pound child. It is to be administered in three equal doses. The drug is available in an oral suspension containing 125 mg/ml. How many mL should be administered per dose?

Ans - 40 pounds = 18.18 kg

18.18 x 5 = 90.9 mg in three divided doses ie 30.3 mg per dose

Oral suspension contains 125 mg / ml

Amount to be administered = Desired dose/Dose on hand x Quantity

ie 30.3/125 x 1 = 0.24 ml

Answer = 0.24ml

2 - The order is for 1.2 million units of penicillin G (Bicillin) IM. Available in 600,000 units/ml. How much should the nurse administer?

Ans - 1 million = 10 lacks

1.2 million = 12 lacks

Available dose = 6 lacks units / ml

Answer =   2 ml dosage has to be administered

3 - The Order is for 2500 ml D5W over 24 hours. Drop factor is 15 gtt/ml. At what rate do you run the IV?

2500 / 24 = 104.17

ie 104.17 ml/ hr

104.17 x 15 /60 ie 1562.55 / 60 = 26.04 gtt / min

Answer = 24.04 gtt / min

4 - Order is for 2000 ml D5W over 24 hours. Drop factor is 60 gtt/ml. What rate do you run the IV?

2000/24 = 83.33

83.33 x 60 / 60 = 83.33 gtt / min

Answer = 83.33 gtt / min

5 - Enoxaparin sodium (Lovenox) 25 mg SC q 12 hours is ordered. The label reads 30 mg/0.3 ml. How much should the nurse administer?

0.3 ml = 30 mg

0.1 ml = 10 mg

So 0.25 ml = 0.25 mg

Answer = 0.25 ml

6 - The order is for meperidine 50 mg IM q 4 hours prn. The label reads meperidine 75 mg/ml. How much should the nurse administer?

ie 50 / 75 x 1 = 0.67 ml

Answer = 0.67 ml per dose

7 - An adult is on continuous IV heparin therapy for thrombophlebitis. The IV contains 15,000 units of heparin in 500 ml of 5% dextrose (D5W) at the rate of 20 ml per hr. How many units per hour is the client receiving?

15000/ 500 = 30 units per ml

30 x 20 = 600 units

Answer = 600 units per hour

8 - The order is for Ancef 1 gram IV in 50 cc 5% dextrose (D5W) to run in over 30 minutes every 6 hours. The administration set delivers 10 gtt/cc. What should the drip rate be

50 x 10 / 30 = 16.67 gtt / min.

Answer = 16.67 gtt / min

9 - The HCP prescribes haloperidol 4 mg tid oral liquid for the patient diagnosed with schizophrenia. How many milliliters should the nurse administer from a bottle labeled “haloperidol 2 mg/ml,” for each dose?

1 ml = 2 mg

Required dosage = 4 mg per dose

Answer = 2 mg